# Frank–Wolfe algorithm for DC optimization problem

### v2

@@ 5.1 Iteration-complexity analysis: The piecewise weak-star-convex case
The recursion (37) together with (39) satisfies the hypotheses of Lemma 2.8, yielding
$f(x^{k})-f_{i_{0}}^{*}\leq\frac{4(L+L_{0})\,\operatorname{diam}({\cal C})^{2}}{k},\qquad\forall k\geq 1,$
and
$\min_{\ell\in\{\lfloor k/2\rfloor+2,\ldots,k\}}|\omega_{\ell}|\leq\frac{16(L+L_{0})\,\operatorname{diam}({\cal C})^{2}}{k-2}.\qquad\forall k\geq 3.$
These are precisely items (i)–(ii) of Theorem 5.5. ∎
Theorem 5.5 shows that, under piecewise–star–convexity, the exact–gradient Frank–Wolfe method with our stepsize achieves the convex-case rate: $f(x_{k})-f^{\star}=\mathcal{O}(1/k)$ and the optimality certificate $|\omega_{k}|$ decays at the same order. In contrast, the next section focuses on the finite–difference variant. We do not pursue complexity bounds there; instead we establish asymptotic convergence under mild consistency assumptions on the difference radius and the adaptive stepsizes: the gradient surrogate converges to $\nabla g$ along the iterates, the scheme is well posed, and every accumulation point is first-order stationary (equivalently, the Frank–Wolfe gap and $|\omega_{k}|$ vanish). Thus the Frank–Wolfe method provides projection-free, asymptotic guarantees compatible with the piecewise–star–convex framework, even without claiming the convex-case $\mathcal{O}(1/k)$ rate.
6 Frank–Wolfe algorithm with finite difference
In this section we introduce a projection-free Frank–Wolfe variant that replaces $\nabla g$ by a finite-difference surrogate with an adaptive stepsize, while retaining the linear minimization oracle of Algorithm 1. The scheme is designed to the DC model f=ghf=g-h from earlier sections, $g$ smooth convex and $h$ convex (via subgradients), and fits the DC piecewise–star–convexity framework developed in previous sections. Our analysis targets asymptotic convergence of the finite-difference method: under mild consistency conditions on the probing radii and stepsizes, the surrogate tracks $\nabla g$, the algorithm is well posed, and every accumulation point is first-order stationary (the Frank–Wolfe gap and $|\omega_{k}|$ vanish). We do not claim rates for this variant; Theorem 5.5 remains the $\mathcal{O}(1/k)$ benchmark for the exact-gradient case. Since piecewise–star–convexity objectives are convex within cells, stabilization in a cell yields local convex geometry, which motivates our stepsize choice.
Let $\{e_{i}\}_{i=1}^{n}$ denote the canonical basis of $\mathbb{R}^{n}$. For a differentiable $g:\mathbb{R}^{n}\to\mathbb{R}$ and a stepsize $s>0$, define the (componentwise) forward finite-difference approximation of $\nabla g(x)$ by
$D(x,s):=\bigl{(}D_{1}(x,s),\ldots,D_{n}(x,s)\bigr{)},\qquad D_{i}(x,s):=\frac{g(x+se_{i})-g(x)}{s},\quad i=1,\ldots,n.$ (40)
The next lemma (see [38, (8.4), p. 195]) provides an upper bound on the error of approximating $\nabla g(x)$ by $D(x,s)$.
Lemma 6.1.
Assume that $g$ satisfies the assumption (A). Then, there holds
$\|\nabla g(x)-D(x,s)\|\leq\frac{{\sqrt{n}L}}{2}s,\qquad\forall x\in\mathbb{R}^{n},\quad\forall s>0.$
Proof.
Setting $\nabla g(x)=(\partial_{1}g(x),\ldots,\partial_{n}g(x))$ and using (40) we have
$\|\nabla g(x)-D(x,s)\|^{2}\leq\sum_{i=1}^{n}\left(\partial_{i}g(x)-D_{i}(x,s)\right)^{2}.$ (41)
Applying Proposition 2.4 with $v=e_{i}$ and $\lambda=s$, we conclude that $D_{i}(x,s)-\partial_{i}g(x)\leq(Ls)/2$. Thus, (41) implies that $\|\nabla g(x)-D(x,s)\|^{2}\leq n((Ls)/2)^{2}$, which implies the desired inequality. ∎
The statement of the Frank–Wolfe algorithm to solve problem (14) with finite difference approximating the gradient $\nabla g$ is as follows:
Algorithm 2.
Frank–WolfeC,f:=g-h algorithm with finite difference
Step 0.
Select $x^{0},x^{1}\in{{\cal C}}$, $x^{0}\neq x^{1}$ and $L_{1}>0$. Set $k=1$.
Step 1.
Take $u^{k}\in\partial h(x^{k})$. Set $j:=\min\{\ell\in{\mathbb{N}}:~2^{\ell}L_{k}\geq 2L_{1}\}$ and $i=0$.
Step 2.
Set
$s_{ij}=\frac{2L_{1}}{2^{i+j}L_{k}\sqrt{n}}\|x^{k}-x^{k-1}\|.$ (42)
Step 3.
Use an “LO oracle” to compute an optimal solution $p^{ij}$ and the optimal value ${\omega}_{ij}(x_{k})$ as follows
$p^{ij}\in{\rm argmin}_{p\in{{\cal C}}}(D(x^{k},s_{ij})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\quad{\omega}_{ij}(x_{k}):=(D(x^{k},s_{ij})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{ij}-x^{k}).$ (43)
Step 4.
If ${\omega}_{ij}(x_{k})<0$, then set $i_{k}=i$ and go to Step 5. Otherwise, set $i\leftarrow i+1$, and go to Step 2.
Step 5.
Compute the step size $\lambda_{j}\in(0,1]$ as follows
$\lambda_{j}=\mbox{min}\left\{1,\frac{|{\omega}_{i_{k}j}(x_{k})|}{{2^{j}L_{k}}\|p^{i_{k}j}-x^{k}\|^{2}}\right\}:={\rm argmin}_{\lambda\in(0,1]}\left\{-|{\omega}_{i_{k}j}(x_{k})|\lambda+\frac{{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda^{2}\right\}.$ (44)
Step 6.
If
$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}-\frac{2^{j}L_{k}}{4}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2},$ (45)
then set $j_{k}=j$ and go to Step 7. Otherwise, set $j\leftarrow j+1$ and go to Step 2.
Step 7.
Set $\lambda_{k}:=\lambda_{j_{k}}$, $p^{k}:=p^{i_{k}j_{k}}$, define the next iterate $x^{k+1}$ and the next approximation to the Lipschitz constant $L_{k+1}$ as follows
Set $k\leftarrow k+1$, and go to Step 1.
Let us describe the main features of Algorithm 2. First of all, note that due to $x^{0},x^{1}\in{\cal C}$, $p^{k}\in{\cal C}$ and $\lambda_{k}\in(0,1]$ for all $k=0,1,\ldots$, and $\cal C$} be convex set, it follows from inductive arguments and first equality in (46) that $(x^{k})_{k\in{\mathbb{R}}}\subset{\cal C}$. Note that (43) implies that
${\omega}_{ij}(x_{k})\leq D(x^{k},s_{ij})^{T}(p-x^{k}),\qquad p\in{\cal C}.$ (47)
Hence, considering that $x^{k}\in{\cal C}$, letting $p=x^{k}$ in the last inequality we conclude that
${\omega}_{ij}(x_{k})\leq 0.$ (48)
In order to simplify the notations, from now on we will use the following two notations:
$s_{k}:=s_{i_{k}j_{k}},\qquad\qquad{\omega}_{k}:={\omega}_{i_{k}j_{k}}(x_{k}).$ (49)
Proposition 6.2.
Assume that in Step 3 we have ${\omega}_{ij}(x_{k})=0$, for all $i\in{\mathbb{N}}$. Then,
$(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k})\geq 0,\qquad\forall p\in{{\cal C}},$ (50)
i.e., $x^{k}$ is a stationary point to problem (14).
Proof.
If in Step 3 we have ${\omega}_{ij}(x_{k})=0$, for all $i\in{\mathbb{N}}$, then it follows from (47) that
$0\leq(D(x^{k},s_{ij})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\forall i\in{\mathbb{N}},\quad\forall p\in{{\cal C}}.$ (51)
Since (42) implies that $\lim_{i\to+\infty}s_{ij}=0$, it follows from (40) that $\lim_{i\to+\infty}D(x^{k},s_{ij})=\nabla g(x^{k})$. Hence, taking limit in (51) we conclude that (50) holds. ∎
In view of Proposition 6.2, (48) and (49) from now on we assume that ${\omega}_{k}<0$, for all $k\in{\mathbb{N}}$.
6.1 Well-definedness
Herein, we establish that the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is well defined.
Lemma 6.3.
Let $s_{ij}$ be define in (42) and assume that ${2^{j}L_{k}}\geq 2L$. Then, there holds
$\|\nabla g(x^{k})-D(x^{k},s_{ij})\|\leq\frac{{L_{1}}}{2}\|x^{k}-x^{k-1}\|.$
Proof.
By applying Lemma 6.1 with $x=x^{k}$ and $s=s_{ij}$ we obtain that
$\|\nabla g(x^{k})-D(x^{k},s_{ij})\|\leq\frac{{\sqrt{n}L}}{2}s_{ij}$
Therefore, using (42) and considering that ${2^{j}L_{k}}\geq 2L$ the desired inequality follows. ∎
Lemma 6.4.
Let $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq 2(L+L_{1}/2)$. Then, (45) holds and $j_{k}$ is the smallest $j\in{\mathbb{N}}$ satisfying the following two conditions $2^{j}L_{k}\geq 2L_{1}$ and (45). Consequently, the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is well defined and satisfies the following inequality:
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\lambda_{k}-\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2},\qquad\qquad k=1,2,\ldots.$ (52)
Proof.
It follows from Proposition 2.4 that
$g(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq g(x^{k})+\nabla g(x^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}+\frac{L}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}.$ (53)
On the other hand, considering that $u^{k}\in\partial h(x^{k})$ it follows from Proposition (2.5) that
$h(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\geq h(x^{k})+\lambda_{j}\langle u^{k},p^{i_{k}j}-x^{k}\rangle$ (54)
$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\\ (\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}+\frac{L-{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}.$ (55)
To proceed with the proof we need first to prove that for $\lambda_{j}$ defined in (44) we have
$|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}\leq-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}.$ (56)
For that we separately consider two cases: $\lambda_{j}=|{\omega}_{i_{k}j}(x_{k})|/{(2^{j}L_{k}\|p^{i_{k}j}-x^{k}\|^{2})}$ and $\lambda_{j}=1$. In the former, we have $2^{j}L_{k}\|p^{i_{k}j}-x^{k}\|^{2}=|{\omega}_{i_{k}j}(x_{k})|/{\lambda_{j}}$,which substituting into the left hand side of (56) yields an equality. If now $\lambda_{j}=1$, then it follows from (44) that $\lambda_{j}=1\leq|{\omega}_{i_{k}j}(x_{k})|/({2^{j}L_{k}}\|p^{i_{k}j}-x^{k}\|^{2})$ or equivalently ${2^{j}L_{k}}\|p^{i_{k}j}-x^{k}\|^{2}\leq|{\omega}_{i_{k}j}(x_{k})|$ and $\lambda_{j}=1$, which again substituting into the left hand side of (56) yields the inequality. Hence, (56) holds. Therefore, (55) becomes
$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\\ (\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}+\frac{L-{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}.$ (57)
Since ${2^{j}L_{k}}\geq 2(L+L_{1}/2)>L$, we can apply Lemma 6.3 to obtain that
$(\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}\leq\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|\|p^{i_{k}j}-x^{k}\|\lambda_{j}.$
Hence, considering that $2\|x^{k}-x^{k-1}\|\|p^{i_{k}j}-x^{k}\|\lambda_{j}\leq\|x^{k}-x^{k-1}\|^{2}+\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}$, we have
$(\nabla g(x^{k})-D(x^{k},s_{i_{k}j}))^{{\scriptscriptstyle\mathrm{T}}}(p^{i_{k}j}-x^{k})\lambda_{j}\leq\frac{L_{1}}{4}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2}.$
Thus, combining the last inequality with (57) and considering that $\lambda_{j}<1$, we conclude that
$f(x^{k}+\lambda_{j}(p^{i_{k}j}-x^{k}))\leq f(x^{k})-\frac{1}{2}|{\omega}_{i_{k}j}(x_{k})|\lambda_{j}+\frac{L+L_{1}/2-{2^{j}L_{k}}}{2}\|p^{i_{k}j}-x^{k}\|^{2}\lambda_{j}^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2}.$
Considering that ${2^{j}L_{k}}\geq 2(L+L_{1}/2)$, the last inequality implies that (45) holds. Since the first trial $j$ in (45) is defined in Step 1, i.e., $j:=\min\{\ell\in{\mathbb{N}}:~2^{\ell}L_{k}\geq 2L_{1}\}$ and $2^{j+1}L_{k}\geq 2^{j}L_{k}$ we conclude that $j_{k}$ is the smallest $j\in{\mathbb{N}}$ satisfying $2^{j}L_{k}\geq 2L_{1}$ and (45). Consequently, $(x^{k})_{k\in{\mathbb{N}}}$ is well defined, which proof the first three statements of the lemma. Therefore, (52) follows from the definition of $j_{k}$ in Step 6, (45) and (46), which concludes the proof. ∎
Since the proof of the next lemma is similar to Lemma 5.2 it will be omitted.
Lemma 6.5.
The sequence $(L_{k})_{k\in{\mathbb{N}}^{*}}$ satisfies the following inequalities
$L_{1}\leq L_{k}\leq 2(L+L_{1}),\quad\qquad k=1,2,\ldots.$ (58)
and the step size sequence $(\lambda_{k})_{k\in{\mathbb{N}}^{*}}$ satisfies
$\lambda_{k}\geq\min\left\{1,\frac{|{\omega}_{k}|}{2(L+L_{1})\operatorname{diam}({{\cal C}})^{2}}\right\},\quad\qquad k=1,2,\ldots.$ (59)
6.2 Convergence analysis
In this section we analyze the convergence properties of the sequence $(x^{k})_{k\in{\mathbb{N}}}$. For that, we assume that the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is infinite.
Lemma 6.6.
The sequence $\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)_{k\in{\mathbb{N}}}$ is monotone and non-increasing. Moreover, there holds:
$\lim_{k\to+\infty}\|x^{k}-x^{k-1}\|=0.$ (60)
As a consequence, $\lim_{k\to+\infty}f(x^{k})={\bar{f}}$ for some ${\bar{f}}\geq f^{*}$.
Proof.
Since $L_{k}\geq L_{1}$, for all $k\in{\mathbb{N}}$, it follows from Lemma 6.4 that
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\lambda_{k}-\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2},\qquad\qquad k=1,2,\ldots.$
which, taking into account that $\frac{1}{2}|{\omega}_{k}|\lambda_{k}\geq 0$, yields
$f(x^{k+1})+\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}\leq f(x^{k})+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2},\qquad\qquad k=1,2,\ldots..$
which implies the first statement of the lemma. We proceed to prove the second statement. For that we first note that Lemma 6.5 implies that $L_{1}\leq L_{k}$, for all $k\in{\mathbb{N}}$. Thus, using again that $\frac{1}{2}|{\omega}_{k}|\lambda_{k}\geq 0$, it follows from (52) that
$\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}-\frac{L_{k}}{4}\|x^{k}-x^{k-1}\|^{2}\leq f(x^{k})-f(x^{k+1}),\qquad k=1,2,\ldots.$ (61)
Let $\ell\in{\mathbb{N}}$ be such that $\ell>1$. Summing up the inequality (61) from $1$ to $\ell$ we obtain
$\sum_{k=1}^{\ell}\Big{(}\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}-\frac{L_{k}}{4}\|x^{k}-x^{k-1}\|^{2}\Big{)}\leq f(x^{1})-f^{*}.$ (62)
Considering that $L_{1}\leq L_{k}$, for all $k\in{\mathbb{N}}$, some algebraic manipulations show that
$\frac{L_{1}}{4}\|x^{1}-x^{0}\|^{2}+\frac{L_{1}}{4}\sum_{k=1}^{\ell}\|x^{k+1}-x^{k}\|^{2}\leq\sum_{k=1}^{\ell}\Big{(}\frac{L_{k+1}}{2}\|x^{k+1}-x^{k}\|^{2}-\frac{L_{k}}{4}\|x^{k}-x^{k-1}\|^{2}\Big{)}.$
Combining (62) with the last inequality we obtain that
$\sum_{k=1}^{\ell}\|x^{k+1}-x^{k}\|^{2}\leq\frac{4}{L_{1}}\Big{(}f(x^{1})-f^{*}+\frac{L_{1}}{4}\|x^{1}-x^{0}\|^{2}\Big{)}.$
Since the last inequality holds for all $\ell\in{\mathbb{N}}$, we have $\sum_{k=1}^{+\infty}\|x^{k+1}-x^{k}\|^{2}\leq+\infty$. Thus, (60) holds. To prove the last statement, we first note that due to $\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)_{k\in\mathbb{N}}$ be a non-increasing sequence and bounded from below by $f^{*}$, it must be convergent. Thus, we set ${\bar{f}}=\lim_{k\to+\infty}\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)$. Hence, considering that
$f(x^{k})=f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}-({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\geq f^{*},$
by using (60) we conclude that $\lim_{k\to+\infty}f(x^{k})={\bar{f}}\geq f^{*}$ and the proof is concluded. ∎
For simplifying the notations we define the following constants
${\bar{\alpha}}:=2(L+L_{1})\operatorname{diam}({{\cal C}})^{2}>0.$ (63)
Lemma 6.7.
The following inequality holds:
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}-\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2},$ (64)
for all $k=1,2,\ldots$.
Proof.
Combining inequalities (52) with (58) and (59), and also using (63), we conclude that
$f(x^{k+1})\leq f(x^{k})-\frac{1}{2}|{\omega}_{k}|\min\left\{1,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|\right\}-\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}+\frac{L_{1}}{4}\|x^{k}-x^{k-1}\|^{2},$
for all $k=1,2,\ldots$. Thus, (64) is an immediate consequence of the previous inequality. ∎
In the next theorem we will use the following constant
$\Gamma:=2(f(x^{1})-f^{*})+{L_{1}}\|x^{2}-x^{1}\|^{2}.$ (65)
Theorem 6.8.
The following statements hold:
  • (i)
    $\lim_{k\to+\infty}{\omega}_{k}=0$. Consequently, every limit point of the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 2 is a stationary point to problem (14);
  • (ii)
    for every $N\in{\mathbb{N}}$, there holds
    $\min\left\{|{\omega}_{k}|:~k=1,\ldots,N\right\}\leq\max\left\{{\Gamma},\sqrt{{\bar{\alpha}}\Gamma}\right\}\frac{1}{\sqrt{N}}.$
Proof.
To prove the item (i), we first note that (64) is equivalent to
$\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}\leq\Big{(}f(x^{k})+\frac{L_{1}}{2}\|x^{k}-x^{k-1}\|^{2}\Big{)}-\Big{(}f(x^{k+1})+\frac{L_{1}}{2}\|x^{k+1}-x^{k}\|^{2}\Big{)},$ (66)
for all $k=1,2,\ldots$. It follows from Lemma 6.6 that $\left(f(x^{k})+({L_{1}}/{2})\|x^{k}-x^{k-1}\|^{2}\right)_{k\in{\mathbb{N}}^{*}}$ monotone and non-increasing. Thus, considering that it is also bounded from below by $f^{*}$, we conclude that it converges. Hence, taking limit in (66) we have
$\lim_{k\to+\infty}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}=0,$
which implies the first statement of item (i). For proving the second statement of item (i), take $\bar{x}$} a limit point of the sequence $(x^{k})_{k\in{\mathbb{N}}}$ and subsequence $(x^{k_{i}})_{i\in{\mathbb{N}}}$ such that $\lim_{i\to+\infty}x^{k_{i}}={\bar{x}}$. Since $\cal C$}} is a compact set the sequence $(u^{k})_{k\in{\mathbb{N}}^{*}}$ is bounded. Thus, without loss of generality we assume that $\lim_{i\to+\infty}u^{k_{i}}={\bar{u}}$. On the other hand, Lemma 6.6 together with (42) imply that
$\lim_{k\to+\infty}s_{k}=0.$ (67)
Let $(s_{k_{i}})_{i\in{\mathbb{N}}^{*}}$ subsequence of $(s_{k})_{k\in{\mathbb{N}}^{*}}$. Thus, we conclude from (40) that
$\lim_{i\to+\infty}D(x^{k_{i}},s_{k_{i}})=\nabla g({\bar{x}}).$ (68)
Finally, using (43) we conclude that
$(D(x^{k_{i}},s_{k_{i}})-u^{k_{i}})^{T}(p-x^{k_{i}})\geq{\omega}_{k_{i}},\qquad\forall p\in{{\cal C}}.$
Therefore, taking limite in the last inequality and using (67) together (68) we obtain that
$(\nabla g({\bar{x}})-{\bar{u}})^{T}(p-{\bar{x}})\geq 0,\qquad{\bar{u}}\in\partial g({\bar{x}}),\qquad\forall~x\in{{\cal C}},$
which implies that $\bar{x}$} is a stationary point to problem (14). To prove the item (ii), take $N\in{\mathbb{N}}^{*}$. Thus, (66) implies that
$\sum_{k=1}^{N}\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}\leq\Big{(}f(x^{1})+\frac{L_{1}}{2}\|x^{2}-x^{1}\|^{2}\Big{)}-\Big{(}f(x^{N+1})+\frac{L_{1}}{2}\|x^{N}-x^{N-1}\|^{2}\Big{)}.$
Hence, we conclude that
$\min\left\{\frac{1}{2}\min\left\{|{\omega}_{k}|,\frac{1}{{\bar{\alpha}}}|{\omega}_{k}|^{2}\right\}:~k=1,\ldots N\right\}\leq\frac{1}{N}\left(f(x^{1})-f^{*}+\frac{L_{1}}{2}\|x^{2}-x^{1}\|^{2}\right).$
It follows from the last inequality that there exists ${\bar{k}}\in\{1,\ldots N\}$ such that
$|{\omega}_{\bar{k}}|\leq\frac{1}{N}\left(2(f(x^{1})-f^{*})+{L_{1}}\|x^{2}-x^{1}\|^{2}\right),\qquad|{\omega}_{\bar{k}}|\leq\sqrt{\frac{{\bar{\alpha}}}{N}\left(2(f(x^{1})-f^{*})+{L_{1}}\|x^{2}-x^{1}\|^{2}\right)},$
which, by using (65), implies the item (ii). ∎

### v3

@@ 5.2 Iteration-complexity analysis
Together with (37), we have $0\leq a_{m}\leq b_{m}$ for all ${m}\geq 1$. Therefore the hypotheses of Lemma 2.8 apply, yielding
$a_{m}\leq\frac{4A}{{m}},\qquad\forall m\geq 1,\qquad\qquad\min_{\tau\in\{\lfloor{m}/2\rfloor+2,\ldots,{m}\}}b_{\tau}\leq\frac{16A}{{m}-2}\qquad\forall m\geq 3.$
Returning to k=N+m1k=N+{m}-1 and considering (38), we conclude that for every $k\geq N$, there holds
$f(x^{k})-f_{i_{*}}^{*}=a_{k-N+1}\leq\frac{4A}{k-N+1},$
and for every $k\geq N+2$, there holds
$\min_{\ell\in\{\lfloor(k-N+1)/2\rfloor+N+1,\ldots,k\}}|\omega_{\ell}|=\min_{\tau\in\{\lfloor(k-N+1)/2\rfloor+2,\ldots,k-N+1\}}b_{\tau}\leq\frac{16A}{k-N-1}.$
Since $A=(L+L_{0})\operatorname{diam}({\cal C})^{2}$, these are exactly the bounds in items (i)(ii). ∎
The next remark rewrites the estimates of Theorem 5.5 in an explicit $\varepsilon$–complexity form. In particular, it provides concrete iteration thresholds ensuring $\varepsilon$–accuracy for the objective residual, with respect to the limiting cell minimum, and for the Frank–Wolfe optimality certificate.
Remark 4.
Under the assumptions of Theorem 5.5, there exist $i_{*}\in\{1,\ldots,m\}$ and $N\in\mathbb{N}$ such that the following $\varepsilon$-complexity estimates hold. Given $\varepsilon>0$, any index $k$ satisfying
$k\ \geq\ N-1+\frac{4}{\varepsilon}(L+L_{0})\operatorname{diam}({\cal C})^{2}$
guarantees $f(x^{k})-f_{i_{*}}^{*}\ \leq\ \varepsilon.$ On the other hand, given $\varepsilon>0$, any index $k$ satisfying the following condition
$k\ \geq\ N+1+\frac{16}{\varepsilon}(L+L_{0})\operatorname{diam}({\cal C})^{2}$
guarantees the existence of an iterate index
$\ell\in\Bigl\{\bigl\lfloor\tfrac{k-N+1}{2}\bigr\rfloor+N+1,\ldots,k\Bigr\}\qquad\text{such that}\qquad|\omega_{\ell}|\ \leq\ \varepsilon.$
Equivalently, after $k$ iterations one can extract, from the last $\mathcal{O}(k)$ iterates, an $\varepsilon$-certificate for the Frank–Wolfe gap. In practice, we do not attempt to estimate $N$. Instead, we use the computable Frank–Wolfe gap as the stopping criterion and terminate as soon as $|\omega_{k}|\leq\varepsilon$ or, more robustly, when $\min_{0\leq t\leq k}|\omega_{t}|\leq\varepsilon$. Theorem 5.5 then guarantees that such a threshold will be reached after a finite number of iterations and provides an $\mathcal{O}(1/\varepsilon)$ worst–case bound, up to the transient index shift encoded by $N$.
To orient the reader, we record a simple situation in which the complexity bounds of Theorem 5.5 apply from the first iterate, that is, one may take $N=0$.
Remark 5.
If all cellwise minimum values coincide, i.e., $f_{i}^{*}=f_{j}^{*}$ for every $i,j$, then in Theorem 5.5 one may take $N=0$. Indeed, no cell has a strictly larger minimum value, hence the index set $\hat{I}=\{\,i:\ f_{i}^{*}>f_{i_{*}}^{*}\,\}$ introduced in Lemma 5.4 is empty. Therefore, the exclusion step is not needed, and the comparison argument leading to (36) holds for every $k$ without any initial index shift. Consequently, the complexity bounds in items (i)–(ii) apply from the first iterate. In particular, this situation covers the globally star–convex case on $\cal C$}, a single–cell partition, for which the unique cell is $\cal C$} itself and $f_{1}^{*}=f^{*}_{\cal C}$. Hence, Theorem 5.5 reduces to the standard convex–type $\mathcal{O}(1/k)$ guarantees for the objective residual $f(x^{k})-f^{*}_{\cal C}$ and for the gap certificate $|\omega_{k}|$.
Theorem 5.5 shows that, under piecewise–star–convexity, the Frank–Wolfe method equipped with our backtracking stepsize achieves the convex-case rate, namely, $f(x^{k})-f^{\star}=\mathcal{O}(1/k)$, and the optimality certificate $|\omega_{k}|$ decays at the same order. Thus, the proposed projection-free scheme enjoys the standard sublinear complexity guarantees while remaining compatible with the piecewise–star–convex framework.